UT 20131214 LUCI2 Nighttime Commissioning

Observer: JHeidt, APramskiy, DThompson
Telescope Operator: GBechetti
LBTO Support: TSargent, DMiller

Plan for the night

- If seeing is good: AGw-transform
- Test new telescope service
- LS-acq. tests
- MOS-acq-tests

Summary

- lost 1/3 night due to high humidity

- did extensive LS-acq. tests, several bugs detected

- DMiller prepared for tomorrows AGw-transform measurement on both sides

Preparations

Alexander & Tom arrived safely.
Daily SkypeCon with Walter at 22.30
Ingo provided a revised MOS-acq tool
Michelle prepared a wide variety of scripts for the upcoming nights, thx smile
Jaron Kurk provided material and scripts for MOS-tests

Details

All times are given in UT

Files are luci2.20131213.xxxx

23:30 LUCI-machine gots stuck, Alexander working on the problem,
seems that some LINUX-libraries died...rebooting machine

00:20 Sunset, mostly clear outside, T_ambient = -5.5deg, T_mirror = 0.5deg,
hum = 95%, wind around 10m/s, dome open to allow for cooling of mirrors

00:30 Humidity went up, need to close, ie no flats. Alexander solved the problem
with the LUCi-machine, problem was with new/old C++-libraries Alexander needs

00:35 0001 0031 ?? Possibly LIR data for readnoise

00:57 Dave started script to determine RON in LIR/MER-modes,

UT01:53
Alexander had to restart some stuff Still closed for humidity

Commit all 14 items in DC_MER.xml script
0093 0094 1x6.8515s
0095 0096 1x6.75s (will also be 6.8515s)
0097 0098 1x7.00s
0099 0100 1x8.00s

0101 0102 1x10.0s
0103 0104 1x12.0s
0105 0106 1x15.0s
0107 0111 1x20.0s
0112 0116 1x60.0s
0117 0121 1x180.0s
0122 0126 1x300.0s
Script paused after end of the 300s sets.

02:00 Humidity has dropped below 90%, we are going to open now,
problems with latch pin(s) for left shutter doors in dome, need to be fixed,
left door still not completely open, may be subject to strong wind

02:43 Geno checked pointing on both sides, did collimation, humidity dropped further,
seeing around 1", so wait a bit with the AGw-transform

LS-acqusition

03:09 Preset to BS9107, didn't work, both LUCI's are authorized, authorize LUCI2 and resend preset, OK
weired readout, #127/128, resend preset, #129-131, OK, problem: ALADIN can not load images,
indicated as filenames are luci2.20131213 without extension! Yak, Ingos new version is not yet
deployed. Did it including the new telescope service GUI, need to recover FW2 inbetween

03:47 Preset to BS9107 again, #133/134, OK, did alignment for one star, looks OK
offsets are not sent from RTD to telescope service => Ingo uses new variable which
is not handled over to telescope service => Alexander tries to fix that

04:24 In the meantime started BS9107, ZP's in JHKKs for N3.75 & N1.8 camera, AM=1.02, #138-177
seeing is variable and degrading a bit 1.2-2", JHKKs-images N1.8 are out of focus
if there is a script which executes obsitems using both cameras, there is an error message by
ALADIN that it has detected 2 different pixel scale values popping up once an image has been taken

AI IStilz has to fix that

04:52 Focus N1.8 camera for reference-filter (K) set back to -325

04:54 Problem with offsets not sent fixed by Alex, preset to BS9107 again, #178/179,
aligned, offset is sent now, #181 verification, offset-sign in x is wrong

05:34 redo to verify wrong sign, Alexander changes sign in x/RA

06:05 Stop, restart SW, need to initialize at IM, YAK, did it also for the MOS where slit is in turnout,
stopped and tried to contact Peter

06:24 OK, Peter fixed it, put mask back to cabinet (we could have also initialized it)

06:32 Some test images using new telescope service GUI, #189-192

06:45 Sign change didn't help, #194/195

06:50 Humidity went up again, closed dome, T=-7deg

07:00 Cooling system for all instruments has problems DAshby needs to come up, alarm went on a few times in the last hours

07:15 Dave & Alexander did some tests: sign calculated for LS-acq. in Ingos tool
in x is wrong (crazy, was finally OK last com), Alexander removed -sign in x in both RTD and telescope manager

07:35 After stop/start SW (in case of a new version) FW2 can not be initialized and need to be recovered

09:20 Humidity dropped, reopened

09:40 Preset BS9126, check on LS-alignmant at PA=0, #195/196, aligned,
confirmation image #197, do not understand this, offset in x has wrong sign (although given properly to TCS)
, and moved only half of the amount requested, offset in y has right sign but is factor 1.5 too large
Position on sky-subracted image: (#196-#195): Source 1150/894, Move-to-Position 988, 1042,
Moved-to-Position 1233/1132 (#197)

10:00 Preset BS9126, check on LS-alignmant without sky-subtraction. #203: Source 1146/1151,
Move-to-position 1231/1232, #204 verification image, moved-to-position 1058/1230
=> Offset in y is OK, in x is reversed in sign. Did some more tests: Offsets from IRAF, RTD
or scripts are OK.

AI TSargent sign-change in telescope service required

Could correctly be done via PCS/GCS but requires some effort, DThompson will talk to JHill about that

10:20 Preset BS9126, check on LS-alignmant with sky-subtraction. #205 (sky), #206 (target).
skysubtracted (#206-#205): Source 1145/1151, move-to-position 1260,1251,
#204 verification image, moved-to-position 947,1162, again wrong offsets (not only in sign in x)
=> suspect that its due to sky-subraction, header in sky-subracted image
contains WCS of background image and not source image

AI IStilz revise tool accordingly

AI TSargent did sign-change in telescope service

11:00 Handover to DMiller for AGw-transform measurement prep for tomorrow on both sides

11:35 Geno went up to the treehouse to reset the rotator (some power glitch)

13:43 DMiller is done with his tests - 25mins before sunrise
Cirrus outside, no flats

13:45 Closing

13:59 Dave started a script to determine persistence using the new "diagonal mask",
#208-210 are 2.75sec darks (test-exps), the rest is persistence determination #211-272

-- JochenHeidt - 14 Dec2013
Topic revision: r11 - 15 Dec 2013, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback